A Solve the initial value problem: y = b[-2 ]y yo) = (^3] B. Find the general solution: I 3 4 y = 1-1 71
Answer #1

kelution A) y= 1946 1374 = my 1 460) = 31 is the characteristic equation det (A-AI)=0 Eigenvedor assouated with A=1 YA-IlX 2CLASSTIME Page No. Date is me characteristic equation det (A-AI) =0 * (3-1) (7-1) +4 =D Egranatue Eigenvertor corresponding t

Please feel free to ask any doubts regarding the solution and please rate positively.

Kind regards.